# Investigating social alignment via mirroring in a system of interacting language models

### v1

@@ Abstract
Figure 1: Illustration of simulated system dynamics with $n=7$ agents with a communication range of $k=3$. Dashed arrows connect agents in the communication range of the receiving agent. Solid arrows connect the agent that interacted with the receiving agent. Arrow color represents the type of interaction – if it is the color of the receiving agent then communicating agent is mirroring, otherwise they are not mirroring. Mirorring occurs with probability $p$ for each interaction. The opinion of an agent and the agents in its range of communication may change at time $t+1$ based off its interaction at time $t$. For example, the blue arrow from the light-blue male to the orange female at $t$ affects the opinion of the orange female at $t+1$ – she becomes light blue and the agents in her communication range change. See Section 1 for further details.
@@ 1 Experimental Design
We let $A_{1},\ldots,A_{n}$ be $n$ agents. Each agent is Meta’s LLaMA-2-7B-Chat (Touvron et al., 2023) equipped with a different external database. As mentioned above, the external databases are a proxy for different knowledge bases and memories of individuals in a community. In our experiment, each database contains sentences about flowers. These sentences are randomly sampled from a collection of sentences originally generated by prompting ChatGPT with “Provide many sentences about various kinds of flowers.".
@@
  • (Measurement Step) At each time $t$, we ask each agent to describe the prettiest flower in a single sentence based on its database. We denote the answer as $B_{i}^{(t)}$ ($i=1,\;\ldots,\;n;\;t=0,\;\ldots,\;T$). The answers will be used to measure agent alignment.
@@
    where $||\cdot||_{2}$ denotes the vector $l_{2}$-norm.
    We note that sentences can be different in many aspects. For example, the sentences “The prettiest flower is rose because roses come in a vast range of colors” and “The prettiest flower is lily because lilies come in a vast range of colors” are different because they are about two types of flowers. The first sentence is also different from “The prettiest flower is rose because its layered petals spiral in a graceful design” because they describe different properties of roses.
    Figure 2: Three patterns of silos: Stable (Pattern I), Unstable (Pattern II), and Decaying (Pattern III). System classification occurs at $T=80$. Top. Example systems for the three of the observed types of patterns. Each line represents an agent. Middle. The dynamics of the number of agents in each silo for each of the example systems in the top row. Each line represents a silo. Bottom. The metrics used to classify a system for each of the example systems in the top row. Systems with stable silos have multiple non-empty silos with unchanging members; systems with unstable silos have multiple non-empty silos with interchanging members; systems with decaying silos have recently experienced the observed minimum entropy.
@@
An illustration of system dynamics is provided in Figure 1. Note that the evolution of the described system depends on the probability of interacting with a mirroring agent ($p$), and the number of nearest neighbors an agent is able to interact with ($k$; sometimes referred to as the agents’ “range of communication"). We measure system evolution by clustering the agents’ responses based on the which flower it describes as the prettiest – we are agnostic to differences in the reason for choosing the flower. In particular, we assign each flower a Flower ID and use it to designate an agent’s cluster (or “silo") membership. Silo membership can change as a function of $t$.
@@ 2.2 The effects of p and k
The splintering of high-$p$, high-$k$ populations is primarily due to the increase in time which a high-$p$ population requires to converge. In particular, when the majority of interactions are instances of a mirroring interaction, there is little opportunity for actual information to be shared. Thus, agents are more likely to remain isolated in their own perspective bubble for a longer amount of time – with only the off-chance that basic interactions serving to collapse perspective bubbles into multi-agent silos – resulting in few instances where a concensus across the entire population is reached by $T=80$. We note that when $p=1$, the number of silos is equal to the number of flower IDs present in the initial conditions, as the opportunity for an informative interaction has been completely replaced by mirroring interactions.
@@ 3 Discussion
Similarly, Carlson and Knoester (2011) and Cinelli et al. (2021) discuss potential causes of the formation of physical and online echo chambers. Essentially, members in family-sized groups reinforce the opinions of fellow members. For physical groups, opinion is fortified based on pre-existing geographical proximity Carlson and Knoester (2011). For online echo chambers, members self-select based on their pre-existing opinions and thus are only ever subjected already-aligned people Cinelli et al. (2021). Different parameter settings of our simulated system allow us to understand the effect of these behaviors when applied to every agent in the system – small $k$ models opinions reinforced by geographical communication constrains and large $p$ models opinions reinforced after self-selecting for interactions with agents with similar opinions. As seen in Figures 4 & 5 these settings lead to many small stable groups that are unable to achieve population consensus due to their isolation from alternative perspectives.
@@ 4 Future Work
For example, we limited our analysis to systems with $n=30$ agents. While $30$ agents is large enough for relatively complicated group dynamics to emerge, simulating larger systems will allow us to draw parallels to the dynamics observed in human systems more easily. Similarly, we focused on system classification at $T=80$. The classification of a system will depend on the choice of $T$, as seen in Figure 6 where we observe unstable silos at $t=80$ and a single silo for $t>110$. Extending our work to longitudinal analysis (i.e., across various $T$) of the systems may be necessary to fully understand group dynamics.
@@
Acknowledgments
We would like to thank Avanti Athreya, Brandon Duderstadt, Henry Farrell, Vince Lyzinski, and Youngser Park for their helpful feedback and discussions throughout the development of this manuscript.

### v2

@@ Abstract
Figure 1: Illustration of simulated system dynamics with $n=7$ agents with a communication range of $k=3$. Dashed arrows connect agents in the communication range of the receiving agent. Solid arrows connect the agent that interacted with the receiving agent. Arrow color represents the type of interaction – if it is the color of the receiving agent then communicating agent is mirroring, otherwise they are not mirroring. Mirroring occurs with probability $p$ for each interaction. The opinion of an agent and the agents in its range of communication may change at time $t+1$ based off its interaction at time $t$. For example, the blue arrow from the light-blue male to the orange female at $t$ affects the opinion of the orange female at $t+1$ – she becomes light blue and the agents in her communication range change. See Section 1 for further details.
@@ 1 Experimental Design
We let $A_{1},\ldots,A_{n}$ be $n$ agents. Each agent is Meta’s LLaMA-2-7B-Chat (Touvron et al., 2023) equipped with a different external database. As mentioned above, the external databases are a proxy for different knowledge bases and memories of individuals in a community. In our experiment, each database contains a fixed number of sentences about flowers. These sentences are randomly sampled from a collection of sentences originally generated by prompting ChatGPT with “Provide sentences that describe the beauty of various kinds of flowers. Sentences should not exceed 25 words.” By limiting the size of database and the length of sentences, we control the initialization consistency across different agents. In our experiment each agent has a unique initialization to properly represent natural diversity in a community.
@@
  • (Measurement Step) At each time $t$, we ask each agent “Describe the prettiest flower in a single sentence.”. We denote the answer as $B_{i}^{(t)}$ ($i=1,\;\ldots,\;n;\;t=0,\;\ldots,\;T$). The answers will be used to measure agent alignment.
@@
    where $||\cdot||_{2}$ denotes the vector $l_{2}$-norm. We use $l_{2}$-norm, as opposed to cosine (dis)similarity, because we expect to extend our framework to the scenario where multiple queries are asked. In the multi-query scenario, $X_{i}^{(t)}$ becomes a matrix instead of a vector, and the Frobenius norm, i.e. $||\cdot||_{2,2}$, is a common choice of metric.
    Figure 2: Three patterns of silos: Stable (Pattern I), Unstable (Pattern II), and Decaying (Pattern III), coming from different pairs of ($p,k$). System classification occurs at $T=80$. Top. Example systems for the three of the observed types of patterns. Each line represents an agent. Middle. The dynamics of the number of agents in each silo for each of the example systems in the top row. Each line represents a silo. Bottom. The metrics used to classify a system for each of the example systems in the top row. Systems with stable silos have multiple non-empty silos with unchanging members; systems with unstable silos have multiple non-empty silos with interchanging members; systems with decaying silos have recently experienced the observed minimum entropy.
@@
An illustration of system dynamics is provided in Figure 1. Note that the evolution of the described system depends on the probability of interacting with a mirroring agent ($p$), and the number of nearest neighbors an agent is able to interact with ($k$; sometimes referred to as the agents’ “range of communication"). We measure system evolution by clustering the agents’ responses based on the flower species that they say is the prettiest. To extract the exact species names from agent responses, we define a comprehensive list of possible flower names and use a Python program to find the names in sentences. For analytical simplicity, we assign each flower a Flower ID and use it to designate an agent’s cluster (or “silo") membership. Silo membership can change as a function of $t$. We note that the embeddings of the descriptions capture the similarities across flower species that we do not fully encode in the Flower IDs.
@@ 2.2 The effects of p and k
The splintering of high-$p$, high-$k$ populations is primarily due to the increase in time which a high-$p$ population requires to converge. In particular, when the majority of interactions are instances of a mirroring interaction, there is little opportunity for actual information to be shared. Thus, agents are more likely to remain isolated in their own perspective bubble for a longer amount of time – with only the off-chance that basic interactions serving to collapse perspective bubbles into multi-agent silos – resulting in few instances where a consensus across the entire population is reached by $T=80$. We note that when $p=1$, the number of silos is equal to the number of flower IDs present in the initial conditions, as the opportunity for an informative interaction has been completely replaced by mirroring interactions.
@@ 3 Discussion
Similarly, Carlson and Knoester (2011) and Cinelli et al. (2021) discuss potential causes of the formation of physical and online echo chambers. Essentially, members in family-sized groups reinforce the opinions of fellow members. For physical groups, opinion is fortified based on pre-existing geographical proximity Carlson and Knoester (2011). For online echo chambers, members self-select based on their pre-existing opinions and thus are only ever subjected already-aligned people Cinelli et al. (2021). Different parameter settings of our simulated system allow us to understand the effect of these behaviors when applied to every agent in the system – small $k$ models opinions reinforced by geographical communication constraints and large $p$ models opinions reinforced after self-selecting for interactions with agents with similar opinions. As seen in Figures 4 & 5 these settings lead to many small stable groups that are unable to achieve population consensus due to their isolation from alternative perspectives.
@@ 4 Limitations
For example, we limited our analysis to systems with $n=30$ agents. While $30$ agents is large enough for relatively complicated group dynamics to emerge, simulating larger systems will allow us to draw parallels to the dynamics observed in human systems more readily. Similarly, we focused on system classification at $T=80$. The classification of a system will depend on the choice of $T$, as seen in Figure 6 where we observe unstable silos at $t=80$ and a single silo for $t>110$. Extending our work to the longitudinal effects (i.e., across various $T$) of system hyperparameters may be necessary to fully understand group dynamics and final consensus behavior.
In addition, our analysis is focused on systems with a fixed ($p,k$) across time. This is a simplification that we make to present digestible results, though we recognize that it is an abstraction that may not properly reflect real-world interactions, as individuals adjust their communication size and level of mirroring based on multiple factors. Extending our analysis to a more dynamic setting may be very valuable.